how to find slope of (4,1) and (8,-5)
Answer by nerdybill(7384)

For any two points on a line:
(x1,y1)
and
(x2,y2)
.
The slope (m) can be found by:
m = (y2 - y1)/(x2 - x1)
.
You were given two points:
(x1,y1):(4,1)
(x2,y2):(8,-5)
.
Plugging it in:
m = (y2 - y1)/(x2 - x1)
m = (-5 - 1)/(8 - 4)
m = (-6)/(4)
m = -3/2 (slope)